OpenStudy (anonymous):

Simplify 7 radical 3 + 8 radical 3 - 2 radical 2

Parth (parthkohli):

There's a property of radicals. \(\Large \color{Orange}{\Rightarrow x\sqrt{a} + y\sqrt{a} = (x + y)\sqrt{a} }\)

Parth (parthkohli):

Similarly, \(\Large \color{Orange}{\Rightarrow 7\sqrt{3} + 8\sqrt{3} = (8 + 7)\sqrt{3} = 15\sqrt{3} }\)

Parth (parthkohli):

That just leaves \(15\sqrt{3} - 2\sqrt{2}\).

OpenStudy (anonymous):

thank you got the same answer!

Parth (parthkohli):

:D